OpenStudy (anonymous):

A certain airline requires that the total outside dimensions (length + width + height) of a carry-on bag not exceed 44 inches. Suppose you want to carry on a bag whose length is twice its height. What is the largest volume bag of this shape that you can carry on a flight? (Round your answer to the nearest integer.)

OpenStudy (welshfella):

L=2h V = 2h^2w

OpenStudy (welshfella):

Also 3h+w <= 44

OpenStudy (pawanyadav):

Height. =x Length =2x Then, width=44-3x Volume =(x)(2x)(44-x) For maximum d(V)/dx=0 Then find. X and put in expression of V

OpenStudy (pawanyadav):

Volume=(x)(2x)(44-3x)